OpenStudy (anonymous):

Find the exact value of the indicated trigonometric function of Ɵ. sin Ɵ = -2/9, tan Ɵ > 0 Find sec Ɵ. Explain please I was absent?

OpenStudy (anonymous):

I completely don't know how to do this haha, and my book doesn't explain it that well. Really confused.

OpenStudy (anonymous):

Ok so sin Ɵ = y / r, so Ɵ = (?, -2) x^2 + (-2)^2 = (-9)^2 x^2 + 4 = 81 x^2 = 77 Am i doing this right?